OpenStudy (anonymous):

if 12 roses cost $18.96, what is the cost for 5 roses???

OpenStudy (anonymous):

12/18.96 = x/5

OpenStudy (anonymous):

18.96 / 12 = 1.58 per rose. 1.58 x 5 = $7.9 for 5 roses

OpenStudy (anonymous):

$7.90

OpenStudy (xishem):

\[12r=$18.96\]\[r=$1.58\]\[5r=?=5($1.58)=$7.90\]

OpenStudy (anonymous):

I'd ask what does 1 cost? then mulyiply the csot of one by 5